# A Malliavin-Gamma calculus approach to Score Based Diffusion Generative models for random fields

### v1

@@ 1 Sample a Gaussian random variable $Y^{\mathrm{EM}}_{0}\sim\mathfrak{m}=\mathcal{N}(0,C)$
4.1. Convergence results for finite-dimensional approximations
$\begin{cases}\mathrm{d}Y_{s}^{J}=-\frac{1}{2}\,Y_{s}^{J}\mathrm{d}s+\sigma_{T-t_{k}}^{-1}\,(\mathrm{v}^{\theta^{\star},J}_{T-t_{k}}(Y_{t_{k}}^{J})-e^{\nicefrac{{T-t_{k}}}{{2}}}Y^{J}_{t_{k}})\mathrm{d}s+\Sigma^{J}\,\mathrm{d}B_{s}\quad\forall\,s\in[t_{k},t_{k+1}]\\ Y_{0}^{J}\sim\mathfrak{m}^{J}=\mathcal{N}(0,(\Sigma^{J})^{2})\,,\end{cases}$
H 1.
there exists $\varepsilon>0$ and $\theta^{\star}\in\Theta$ such that
$\frac{1}{M}\,\sum_{j=1}^{M}\,\mathbb{E}\left\|\mathbb{E}[\overrightarrow{X}^{J}_{0}|\overrightarrow{X}^{J}_{T-t_{j}}=\overleftarrow{X}^{J}_{t_{j}}]-\mathrm{v}_{T-t_{j}}^{\theta^{\star},J}(\overrightarrow{X}^{J}_{T-t_{j}})\right\|_{(\Sigma^{J})^{-1}}^{2}\leq\varepsilon^{2}\,,$
$\frac{1}{M}\,\sum_{j=1}^{M}\,\mathbb{E}\left\|\mathbb{E}[\overrightarrow{X}_{0}|\overrightarrow{X}_{T-t_{j}}=\overleftarrow{X}_{t_{j}}]-\mathrm{v}_{T-t_{j}}^{\theta^{\star}}(\overrightarrow{X}_{T-t_{j}})\right\|_{\mathcal{H}_{\mathfrak{m}}}^{2}\leq\varepsilon^{2}\,.$
We are now ready to prove our $\mathrm{KL}$-convergence bound.
Theorem 2.
Proof.
Since $\mu_{\mathrm{data}}^{J}=\mathcal{L}(\overleftarrow{X}_{T}^{J})$, from [20, Theorem 2.3] and from the data-processing inequality [30, Lemma 1.6], we immediately deduce that
(4.3) $\mathrm{KL}(\mu_{\mathrm{data}}^{J}|\mathcal{L}(Y_{T}^{J}))\leq$ $\,\mathrm{KL}(\mathcal{L}(\overleftarrow{X}^{J}_{\cdot})|\mathcal{L}(Y^{J}_{\cdot}))=\mathrm{KL}(\mathcal{L}(\overleftarrow{\mathrm{P}}_{0}^{J})|\mathcal{L}(Y^{J}_{0}))$
$\qquad\qquad+\frac{1}{2}\sum_{k=0}^{M-1}\int_{t_{k}}^{t_{k+1}}\mathbb{E}_{\overleftarrow{\mathrm{P}}^{J}}\left\|V^{J}_{s}(\overleftarrow{X}^{J}_{s})-\mathrm{v}^{\theta^{\star},J}_{T-t_{k}}(\overleftarrow{X}^{J}_{t_{k}})+e^{\nicefrac{{T-t_{k}}}{{2}}}\overleftarrow{X}_{t_{k}}^{J}\right\|_{(\Sigma^{J})^{-1}}^{2}\mathrm{d}s$
$\leq\,\mathrm{KL}$ $(\overrightarrow{\mathrm{P}}_{T}^{J}|\mathfrak{m}^{J})+\sum_{k=0}^{M-1}\int_{t_{k}}^{t_{k+1}}\mathbb{E}_{\overleftarrow{\mathrm{P}}^{J}}\left\|V^{J}_{t_{k}}(\overleftarrow{X}^{J}_{t_{k}})-\mathrm{v}^{\theta^{\star},J}_{T-t_{k}}(\overleftarrow{X}^{J}_{t_{k}})+e^{\nicefrac{{T-t_{k}}}{{2}}}\overleftarrow{X}_{t_{k}}^{J}\right\|_{(\Sigma^{J})^{-1}}^{2}\mathrm{d}s$
$\qquad\qquad\qquad+\sum_{k=0}^{M-1}\int_{t_{k}}^{t_{k+1}}\mathbb{E}_{\overleftarrow{\mathrm{P}}^{J}}\left\|V^{J}_{s}(\overleftarrow{X}^{J}_{s})-V^{J}_{t_{k}}(\overleftarrow{X}^{J}_{t_{k}})\right\|_{(\Sigma^{J})^{-1}}^{2}\mathrm{d}s\,,$
where for notations’ sake we have introduced the (finite-dimensional relative) score function (see for instance [34, Lemma 1])
which combined with the latter gives
$\frac{\mathrm{d}}{\mathrm{d}t}\mathrm{KL}(\overrightarrow{\mathrm{P}}_{t}|\mathfrak{m})=\frac{\mathrm{d}}{\mathrm{d}t}\int(P_{t}\rho_{0})\log(P_{t}\rho_{0})\mathrm{d}\mathfrak{m}=\int(1+\log(P_{t}\rho_{0}))L(P_{t}\rho_{0})\mathrm{d}\mathfrak{m}=-4\int\frac{\Gamma(\rho_{t})}{\rho_{t}}\mathrm{d}\mathfrak{m}$
$=-4\int\left\|\bar{\nabla}\sqrt{\rho_{t}}\right\|_{\mathcal{H}_{\mathfrak{m}}}^{2}\mathrm{d}\mathfrak{m}\leq-\frac{1}{2}\mathrm{KL}(\overrightarrow{\mathrm{P}}_{t}|\mathfrak{m})\,.$
Lastly, let us focus on the last term in (4.3). In order to do that, hereafter let us assume $\mu_{\mathrm{data}}$ to have finite eight moments, this extra assumption can be lifted via a standard approximation procedure as in [10, Lemma 3]. Next, notice that by reasoning as in [10, Proposition 2 and Equation (27)] we can deduce (see Section 4.1) that
and
(4.6) $\mathrm{d}\left\|V^{J}_{t}(\overleftarrow{X}^{J}_{t})\right\|_{(\Sigma^{J})^{-1}}^{2}=\biggl{(}\left\|V^{J}_{t}(\overleftarrow{X}^{J}_{t})\right\|_{(\Sigma^{J})^{-1}}^{2}+\left\|\nabla V^{J}_{t}(\overleftarrow{X}^{J}_{t})\right\|^{2}_{\mathrm{Fr}}\biggr{)}\mathrm{d}t+2\,V^{J}_{t}(\overleftarrow{X}^{J}_{t})\,\cdot(\Sigma^{J})^{-1}\nabla V^{J}_{t}(\overleftarrow{X}_{t}^{J})\mathrm{d}B_{t}\,,$
$\mathbb{E}_{\overleftarrow{\mathrm{P}}^{J}}\left\|V^{J}_{s}(\overleftarrow{X}^{J}_{s})-V^{J}_{t_{k}}(\overleftarrow{X}^{J}_{t_{k}})\right\|_{(\Sigma^{J})^{-1}}^{2}=\mathbb{E}_{\overleftarrow{\mathrm{P}}^{J}}\left\|\frac{1}{2}\int_{t_{k}}^{s}V^{J}_{u}(\overleftarrow{X}_{u}^{J})\mathrm{d}u+\int_{t_{k}}^{s}\Sigma^{J}\nabla V^{J}_{u}(\overleftarrow{X}_{u}^{J})\,\mathrm{d}B_{u}\right\|_{(\Sigma^{J})^{-1}}^{2}$
and hence that the third term in (4.3) can be bounded as follows
$\leq$ $\,h\,\max\{2,\nicefrac{{h}}{{2}}\}\,\mathbb{E}_{\overleftarrow{\mathrm{P}}^{J}}\left\|V^{J}_{T}(\overleftarrow{X}^{J}_{T})\right\|_{(\Sigma^{J})^{-1}}^{2}=h\,\max\{2,\nicefrac{{h}}{{2}}\}\,\mathbb{E}_{\overrightarrow{\mathrm{P}}^{J}}\left\|\Sigma^{J}\nabla\log\rho^{J}_{0}(\overrightarrow{X}_{0}^{J})\right\|^{2}_{2},$
where $[\Phi_{J},\Phi_{J^{\perp}}]_{\mathrm{H}}\in\mathrm{H}$ denotes the element whose coordinates are $(\Phi_{J},\Phi_{J^{\perp}})$ and where we have defined $\Phi^{J^{\perp}}$ as the projection of $\Phi\in\mathrm{H}$ onto the orthogonal complement to $\mathrm{Span}\{Y_{\ell_{j}}\colon j=1,\dots,N\}$, that is the subspace generated by the element of the orthonormal basis not in $J$, whereas $\mathfrak{m}^{J^{\perp}}$ is simply the pushforward under this orthogonal projection of $\mathfrak{m}$. If $e_{j}$ denotes the $j^{th}$ element of the standard basis for $\mathbb{R}^{N}$, for any $j=1,\dots,N$ it holds
$=\frac{\biggl{(}\int\mathrm{d}_{t_{|t=0}}\rho_{0}([\Phi_{J},\Phi_{J^{\perp}}]_{\mathrm{H}}+t\,Y_{\ell_{j}})\mathrm{d}\mathfrak{m}^{J^{\perp}}(\Phi_{J^{\perp}})\biggr{)}^{2}}{\int 4\,\rho_{0}([\Phi_{J},\Phi_{J^{\perp}}]_{\mathrm{H}})\mathrm{d}\mathfrak{m}^{J^{\perp}}(\Phi_{J^{\perp}})}\leq\int\frac{(\mathrm{d}_{t_{|t=0}}\rho_{0}([\Phi_{J},\Phi_{J^{\perp}}]_{\mathrm{H}}+t\,Y_{\ell_{j}})\,)^{2}}{4\,\rho_{0}([\Phi_{J},\Phi_{J^{\perp}}]_{\mathrm{H}})}\mathrm{d}\mathfrak{m}^{J^{\perp}}(\Phi_{J^{\perp}})$
$=\int\biggl{(}\mathrm{d}_{t_{|t=0}}\sqrt{\rho_{0}}([\Phi_{J},\Phi_{J^{\perp}}]_{\mathrm{H}}+t\,Y_{\ell_{j}})\biggr{)}^{2}\mathrm{d}\mathfrak{m}^{J^{\perp}}(\Phi_{J^{\perp}})=\int\langle\bar{\nabla}\sqrt{\rho_{0}}([\Phi_{J},\Phi_{J^{\perp}}]_{\mathrm{H}}),\,Y_{\ell_{j}}\rangle_{\mathcal{H}_{\mathfrak{m}}}^{2}\,\mathrm{d}\mathfrak{m}^{J^{\perp}}(\Phi_{J^{\perp}})\,.$
From this we may finally deduce that
$=4\,\mathbb{E}_{\mathfrak{m}^{J}}\biggl{[}\sum_{j=1}^{N}C_{\ell_{j}}\,(\partial_{j}\sqrt{\rho^{J}_{0}})^{2}\biggr{]}\leq 4\,\mathbb{E}_{\mathfrak{m}^{J}}\biggl{[}\sum_{j=1}^{N}C_{\ell_{j}}\,\int\langle\bar{\nabla}\sqrt{\rho_{0}}([\Phi_{J},\Phi_{J^{\perp}}]_{\mathrm{H}}),\,Y_{\ell_{j}}\rangle_{\mathcal{H}_{\mathfrak{m}}}^{2}\,\mathrm{d}\mathfrak{m}^{J^{\perp}}(\Phi_{J^{\perp}})\biggr{]}$
By combining (4.7) and (4.8) we deduce that
which corresponds to our thesis. ∎
For completeness, below we prove (4.5) and (4.6).
[…]
$\mathrm{d}\left\|V^{J}_{t}(\overleftarrow{X}^{J}_{t})\right\|_{(\Sigma^{J})^{-1}}^{2}=\biggl{(}\left\|V^{J}_{t}(\overleftarrow{X}^{J}_{t})\right\|_{(\Sigma^{J})^{-1}}^{2}+\left\|\nabla V^{J}_{t}(\overleftarrow{X}^{J}_{t})\right\|^{2}_{\mathrm{Fr}}\biggr{)}\mathrm{d}t+2\,V^{J}_{t}(\overleftarrow{X}^{J}_{t})\,\cdot(\Sigma^{J})^{-1}\nabla V^{J}_{t}(\overleftarrow{X}_{t}^{J})\mathrm{d}B_{t}\,,$
Proof.
which implies
5. Generative model for spherical random fields
@@
In this last section we specify our results to the spherical random fields case, i.e., when the state space corresponds to the Hilbert space $\mathrm{H}={\mathrm{L}^{2}({\mathbb{S}^{2}})}$. For a comprehensive introduction to spherical random fields we refer the interested reader to the monograph [26]. For sake of completeness we introduce here the objects that will appear below. A spherical random field can be seen as a random element $T\in{\mathrm{L}^{2}({\mathbb{S}^{2}})}$. A nice characteristic is that it can be decomposed in a spectral decomposition (also known as the Karhunen-Loève decomposition, which holds almost surely thanks to the deterministic Peter-Weyl Theorem) as
@@
$\mathcal{H}_{\mathfrak{m}^{\kappa,\beta}}\coloneqq\biggl{\{}u=\sum_{\ell,m}a_{\ell,m}Y_{\ell,m}\in{\mathrm{L}^{2}({\mathbb{S}^{2}})}\quad\text{ s.t. }\quad\sum_{\ell\geq 0}\sum_{m=-\ell}^{\ell}a_{\ell,m}^{2}(\kappa^{2}+\ell(\ell+1))^{2\beta}<+\infty\biggr{\}}\,,$
@@
$\begin{cases}\mathrm{d}\overrightarrow{a}_{\ell,m}(s)=-\frac{1}{2}\,\overrightarrow{a}_{\ell,m}(s)+(\kappa^{2}+\ell(\ell+1))^{-\beta}\,\mathrm{d}B_{s}\\ \overrightarrow{a}_{\ell,m}(0)=\langle u_{\mathrm{data}},Y_{\ell,m}\rangle_{{\mathrm{L}^{2}({\mathbb{S}^{2}})}}\,,\end{cases}$
where $u_{\mathrm{data}}\sim\mu_{\mathrm{data}}$ is a sample from our data distribution. Once the forward dynamics for the $\overrightarrow{a}_{\ell,m}$ has been simulated, we may train a neural network as in Algorithm 1 in order to learn an approximate score function $\mathrm{v}^{\theta^{\star}}$ and then generate the new data samples $Y^{\mathrm{gen}}_{T}$ where
@@
$\begin{cases}\mathrm{d}g_{\ell,m}(s)=-\frac{1}{2}\,g_{\ell,m}(s)\mathrm{d}s+\sigma_{T-t_{j}}^{-1}\,(\mathrm{v}^{\theta^{\star}}_{T-t_{j}}(Y^{\mathrm{gen}}_{t_{j}})-e^{\nicefrac{{T-t_{j}}}{{2}}}g_{\ell,m}(t_{j}))\mathrm{d}s+(\kappa^{2}+\ell(\ell+1))^{-\beta}\,\mathrm{d}B_{s}\\ \text{and it starts with }g_{\ell,m}(0)\sim\mathcal{N}(0,(\kappa^{2}+\ell(\ell+1))^{-2\beta})\,.\end{cases}$

### v2

@@ 1 Sample a Gaussian random variable $Y^{\mathrm{EM}}_{0}\sim\mathfrak{m}=\mathcal{N}(0,C)$
4.1. Generative model for spherical random fields
@@
In this section we specify our infinite-dimensional generative model to the spherical random fields case, i.e., when the state space corresponds to the Hilbert space $\mathrm{H}={\mathrm{L}^{2}({\mathbb{S}^{2}})}$. For a comprehensive introduction to spherical random fields we refer the interested reader to the monograph [29]. For sake of completeness we introduce here the objects that will appear below. A spherical random field can be seen as a random element $T\in{\mathrm{L}^{2}({\mathbb{S}^{2}})}$. A nice characteristic is that it can be decomposed in a spectral decomposition (also known as the Karhunen-Loève decomposition, which holds almost surely thanks to the deterministic Peter-Weyl Theorem) as
@@
$\mathcal{H}_{\mathfrak{m}^{\kappa,\beta}}\coloneqq\biggl\{u=\sum_{\ell,m}a_{\ell,m}Y_{\ell,m}\in{\mathrm{L}^{2}({\mathbb{S}^{2}})}\;\text{ s.t. }\;\sum_{\ell\geq 0}\sum_{m=-\ell}^{\ell}a_{\ell,m}^{2}(\kappa^{2}+\ell(\ell+1))^{2\beta}<+\infty\biggr\}\,$,
@@
[…]
where $u_{\mathrm{data}}\sim\mu_{\mathrm{data}}$ is a sample from our data distribution.
Once the forward dynamics for the $\overrightarrow{a}_{\ell,m}$ has been simulated, we may train a neural network as in Algorithm 1 in order to learn an approximate score function $\mathrm{v}^{\theta^{\star}}$ and then generate the new data samples $Y^{\mathrm{gen}}_{T}$ where
@@
$\begin{cases}\begin{aligned} \mathrm{d}g_{\ell,m}(s)=-\frac{1}{2}\,g_{\ell,m}(s)\mathrm{d}s+\sigma_{T-t_{j}}^{-1}\,(\mathrm{v}^{\theta^{\star}}_{T-t_{j}}(Y^{\mathrm{gen}}_{t_{j}})-e^{\nicefrac{{T-t_{j}}}{{2}}}g_{\ell,m}(t_{j}))\mathrm{d}s\\ +(\kappa^{2}+\ell(\ell+1))^{-\beta}\,\mathrm{d}B_{s}\end{aligned}\\ \text{and it starts with }g_{\ell,m}(0)\sim\mathcal{N}(0,(\kappa^{2}+\ell(\ell+1))^{-2\beta})\,.\end{cases}$
@@
4.1.1. Numerical simulation
Figure 1. Simulation of the forward noising process at different time-steps.
In Figure 1 we simulate the forward noising dynamics described in our spherical random fields example. In this example, we have considered an ERA5 map from the European Centre for Medium-Range Weather Forecasts [18, 11] plotting the air temperature at 2 meters above the surface of the Earth on the 01/01/2023 (in Kelvin). In order to fit our framework we have firstly turned this ERA5 map into an Healpy map (top left corner in Figure 1), in order to manipulate it with the Python package Healpy useful for spherical random fields simulations, which is based on the HEALPix [^2] http://healpix.sourceforge.net C++ library [46, 15]. Then, we have computed the (real) harmonic coefficients $(a_{\ell,m})_{\ell,m}$. To be more precise, we have firstly applied the function healpy.sphtfunc.map2alm, with $L_{\max}=128$, which returns the complex harmonic coefficients associated to a given Healpy map and then from the latters we have computed the real ones.
Then, we have applied the noising dynamics (OU evolution) to the vector of the (real) harmonic coefficients $(a_{\ell,m})_{\ell,m}$ as described in the previous section, with target noise equal to a Whittle-Matérn field with $\kappa=8$ and $\beta=0.65$ (bottom right corner in Figure 1). For the simulation of the OU process we have explicitly computed the time evolution in $10^{3}$ time-steps equidistributed in $[0,1]$ and in $10^{3}$ steps equidistributed in $[1,20]$.
Finally, from the OU evolution of the $(a_{\ell,m})_{\ell,m}$ we have computed the corresponding evolution for the complex harmonic coefficients and built the time-evolution of the Healpy maps using the function healpy.sphtfunc.alm2map. The plot has been obtained using healpy.visufunc.mollview.
We have stopped our process at $T_{\max}=20$ since at this stage our empirical angular power spectrum $(\hat{C}_{\ell})_{\ell\in[0,128]}$ (i.e., the one computed by averaging over $m$ the harmonic coefficients $(a_{\ell,m})_{m=-\ell,\dots,0,\dots,+\ell}$) fits the theoretical angular power spectrum of the Whittle-Matérn field given at (4.3). In Figure 2 and Figure 3 we plot the theoretical angular power spectrum, the empirical one, and an empirical power spectrum where we average over the last $30$ time-steps of the OU evolution (to compensate the error in the low frequencies of $\hat{C}_{\ell}$, due to the lack of enough harmonic coefficients to average out for low values of $\ell$). Lastly, in Figure 4 we plot the ratio between the empirical (and averaged empirical) $\hat{C}_{\ell}$ and the theoretical $C_{\ell}$.
Figure 2. Plot of empirical and theoretical power spectra
Figure 3. Plot of averaged (over last 30 steps), empirical and theoretical power spectra
Figure 4. Ratio between empirical $\hat{C}_{\ell}$ and theoretical $C_{\ell}$ (on the left) and ratio between averaged empirical $\hat{C}_{\ell}$ and theoretical $C_{\ell}$ (on the right)
Given the above preprocessing of the ERA5 maps into Healpy maps and real vector of harmonic coefficients, the training algorithm described in Algorithm 1 can be obtained by adapting any standard architecture already employed in the literature for vector-valued diffusion models, taking into account the OU dynamics we have choosen. Once this is done, Algorithm 2 is straightforward. Implementing the generative model was out of the scope of the current paper, which, on the other hand, investigates the theoretical point of view of infinite-dimensional-generative models and analyses theoretical convergence bounds (cf. Section 5).
The purpose of the current section is merely visualizing the noising forward dynamics of the harmonic coefficients in terms of spherical fields, in order to help the target readership (unfamiliar with generative models) in understanding what happens along the noising algorithm. In view of that, let us conclude by mentioning that the noising forward procedure (which takes places along Algorithm 1) should be thought as following Figure 1 from the top-left corner towards the bottom-right corner, whereas the denoising procedure (i.e., the sampling algorithm Algorithm 2) can be seen as following Figure 1 from the bottom-right corner towards the top-left corner.
5. Convergence results for finite-dimensional approximations
[…]
H 1.
there exists $\varepsilon>0$ and $\theta^{\star}\in\Theta$ such that
$\frac{1}{M}\,\sum_{j=1}^{M}\,\mathbb{E}\left\|\mathbb{E}[\overrightarrow{X}^{J}_{0}|\overrightarrow{X}^{J}_{T-t_{j}}=\overleftarrow{X}^{J}_{t_{j}}]-\mathrm{v}_{T-t_{j}}^{\theta^{\star},J}(\overleftarrow{X}^{J}_{t_{j}})\right\|_{(\Sigma^{J})^{-1}}^{2}\leq\varepsilon^{2}\,$,
$\frac{1}{M}\,\sum_{j=1}^{M}\,\mathbb{E}\left\|\mathbb{E}[\overrightarrow{X}_{0}|\overrightarrow{X}_{T-t_{j}}=\overleftarrow{X}_{t_{j}}]-\mathrm{v}_{T-t_{j}}^{\theta^{\star}}(\overleftarrow{X}_{t_{j}})\right\|_{\mathcal{H}_{\mathfrak{m}}}^{2}\leq\varepsilon^{2}\,.$
We are now ready to prove our $\mathrm{KL}$-convergence bound.
Theorem 4.
Proof.
Since $\mu_{\mathrm{data}}^{J}=\mathcal{L}(\overleftarrow{X}_{T}^{J})$, from [23, Theorem 2.3] and from the data-processing inequality [35, Lemma 1.6], we immediately deduce that
(5.1) $\mathrm{KL}(\mu_{\mathrm{data}}^{J}|\mathcal{L}(Y_{T}^{J}))\leq\,\mathrm{KL}(\mathcal{L}(\overleftarrow{X}^{J}_{\cdot})|\mathcal{L}(Y^{J}_{\cdot}))$
$=$ $\,\mathrm{KL}(\mathcal{L}(\overleftarrow{\mathrm{P}}_{0}^{J})|\mathcal{L}(Y^{J}_{0}))$
$\quad+\frac{1}{2}\sum_{k=0}^{M-1}\int_{t_{k}}^{t_{k+1}}\mathbb{E}_{\overleftarrow{\mathrm{P}}^{J}}\left\|V^{J}_{s}(\overleftarrow{X}^{J}_{s})-\sigma_{T-t_{k}}^{-1}(\mathrm{v}^{\theta^{\star},J}_{T-t_{k}}(\overleftarrow{X}^{J}_{t_{k}})-e^{\nicefrac{{T-t_{k}}}{{2}}}\overleftarrow{X}_{t_{k}}^{J})\right\|_{(\Sigma^{J})^{-1}}^{2}\mathrm{d}s$
$\leq$ $\,\mathrm{KL}(\overrightarrow{\mathrm{P}}_{T}^{J}|\mathfrak{m}^{J})$
$\quad+\sum_{k=0}^{M-1}\int_{t_{k}}^{t_{k+1}}\mathbb{E}_{\overleftarrow{\mathrm{P}}^{J}}\left\|V^{J}_{t_{k}}(\overleftarrow{X}^{J}_{t_{k}})-\sigma_{T-t_{k}}^{-1}(\mathrm{v}^{\theta^{\star},J}_{T-t_{k}}(\overleftarrow{X}^{J}_{t_{k}})-e^{\nicefrac{{T-t_{k}}}{{2}}}\overleftarrow{X}_{t_{k}}^{J})\right\|_{(\Sigma^{J})^{-1}}^{2}\mathrm{d}s$
$\quad+\sum_{k=0}^{M-1}\int_{t_{k}}^{t_{k+1}}\mathbb{E}_{\overleftarrow{\mathrm{P}}^{J}}\left\|V^{J}_{s}(\overleftarrow{X}^{J}_{s})-V^{J}_{t_{k}}(\overleftarrow{X}^{J}_{t_{k}})\right\|_{(\Sigma^{J})^{-1}}^{2}\mathrm{d}s\,$,
where for notations’ sake we have introduced the (finite-dimensional relative) score function (see for instance [38, Lemma 1])
which combined with the latter gives
$\frac{\mathrm{d}}{\mathrm{d}t}\mathrm{KL}(\overrightarrow{\mathrm{P}}_{t}|\mathfrak{m})=\frac{\mathrm{d}}{\mathrm{d}t}\int(P_{t}\rho_{0})\log(P_{t}\rho_{0})\mathrm{d}\mathfrak{m}=\int(1+\log(P_{t}\rho_{0}))L(P_{t}\rho_{0})\mathrm{d}\mathfrak{m}$
$=-4\int\frac{\Gamma(\rho_{t})}{\rho_{t}}\mathrm{d}\mathfrak{m}=-4\int\left\|\bar{\nabla}\sqrt{\rho_{t}}\right\|_{\mathcal{H}_{\mathfrak{m}}}^{2}\mathrm{d}\mathfrak{m}\leq-\frac{1}{2}\mathrm{KL}(\overrightarrow{\mathrm{P}}_{t}|\mathfrak{m})\,.$
Lastly, let us focus on the last term in (5.1). In order to do that, hereafter let us assume $\mu_{\mathrm{data}}$ to have finite eight moments, this extra assumption can be lifted via a standard approximation procedure as in [10, Lemma 3]. Next, notice that by reasoning as in [10, Proposition 2 and Equation (27)] we can deduce (see Section 5) that
and
(5.4) $\mathrm{d}\left\|V^{J}_{t}(\overleftarrow{X}^{J}_{t})\right\|_{(\Sigma^{J})^{-1}}^{2}=\biggl(\left\|V^{J}_{t}(\overleftarrow{X}^{J}_{t})\right\|_{(\Sigma^{J})^{-1}}^{2}+\left\|\nabla V^{J}_{t}(\overleftarrow{X}^{J}_{t})\right\|^{2}_{\mathrm{Fr}}\biggr)\mathrm{d}t$
$+2\,V^{J}_{t}(\overleftarrow{X}^{J}_{t})\,\cdot(\Sigma^{J})^{-1}\nabla V^{J}_{t}(\overleftarrow{X}_{t}^{J})\mathrm{d}B_{t}\,$,
$=$ $\,\mathbb{E}_{\overleftarrow{\mathrm{P}}^{J}}\left\|\frac{1}{2}\int_{t_{k}}^{s}V^{J}_{u}(\overleftarrow{X}_{u}^{J})\mathrm{d}u+\int_{t_{k}}^{s}\Sigma^{J}\nabla V^{J}_{u}(\overleftarrow{X}_{u}^{J})\,\mathrm{d}B_{u}\right\|_{(\Sigma^{J})^{-1}}^{2}$
and hence that the third term in (5.1) can be bounded as follows
$\leq$ $\,h\,\max\{2,\nicefrac{{h}}{{2}}\}\,\mathbb{E}_{\overleftarrow{\mathrm{P}}^{J}}\left\|V^{J}_{T}(\overleftarrow{X}^{J}_{T})\right\|_{(\Sigma^{J})^{-1}}^{2}$
$=$ $\,h\,\max\{2,\nicefrac{{h}}{{2}}\}\,\mathbb{E}_{\overrightarrow{\mathrm{P}}^{J}}\left\|\Sigma^{J}\nabla\log\rho^{J}_{0}(\overrightarrow{X}_{0}^{J})\right\|^{2}_{2},$
where $\Phi_{J},\Phi_{J^{\perp}}]_{\mathrm{H}}\in\mathrm{H}$ denotes the element whose coordinates are $(\Phi_{J},\Phi_{J^{\perp}})$ and where we have defined $\Phi^{J^{\perp}}$ as the projection of $\Phi\in\mathrm{H}$ onto the orthogonal complement to $\mathrm{Span}\{Y_{\ell_{j}}\colon j=1,\dots,N\}$, that is the subspace generated by the element of the orthonormal basis not in $J$, whereas $\mathfrak{m}^{J^{\perp}}$ is simply the pushforward under this orthogonal projection of $\mathfrak{m}$. If $e_{j}$ denotes the $j^{th}$ element of the standard basis for $\mathbb{R}^{N}$, for any $j=1,\dots,N$ it holds
$=$ $\,\frac{\biggl(\int\mathrm{d}_{t_{|t=0}}\rho_{0}([\Phi_{J},\Phi_{J^{\perp}}]_{\mathrm{H}}+t\,Y_{\ell_{j}})\mathrm{d}\mathfrak{m}^{J^{\perp}}(\Phi_{J^{\perp}})\biggr)^{2}}{\int 4\,\rho_{0}([\Phi_{J},\Phi_{J^{\perp}}]_{\mathrm{H}})\mathrm{d}\mathfrak{m}^{J^{\perp}}(\Phi_{J^{\perp}})}$
$\leq$ $\,\int\frac{(\mathrm{d}_{t_{|t=0}}\rho_{0}([\Phi_{J},\Phi_{J^{\perp}}]_{\mathrm{H}}+t\,Y_{\ell_{j}})\,)^{2}}{4\,\rho_{0}([\Phi_{J},\Phi_{J^{\perp}}]_{\mathrm{H}})}\mathrm{d}\mathfrak{m}^{J^{\perp}}(\Phi_{J^{\perp}})$
$=$ $\,\int\biggl(\mathrm{d}_{t_{|t=0}}\sqrt{\rho_{0}}([\Phi_{J},\Phi_{J^{\perp}}]_{\mathrm{H}}+t\,Y_{\ell_{j}})\biggr)^{2}\mathrm{d}\mathfrak{m}^{J^{\perp}}(\Phi_{J^{\perp}})$
$=$ $\,\int\langle\bar{\nabla}\sqrt{\rho_{0}}([\Phi_{J},\Phi_{J^{\perp}}]_{\mathrm{H}}),\,Y_{\ell_{j}}\rangle_{\mathcal{H}_{\mathfrak{m}}}^{2}\,\mathrm{d}\mathfrak{m}^{J^{\perp}}(\Phi_{J^{\perp}})\,.$
From this we may finally deduce that
$=$  $4\,\mathbb{E}_{\mathfrak{m}^{J}}\biggl[\sum_{j=1}^{N}C_{\ell_{j}}\,(\partial_{j}\sqrt{\rho^{J}_{0}})^{2}\biggr]$
$\leq$  $4\,\mathbb{E}_{\mathfrak{m}^{J}}\biggl[\sum_{j=1}^{N}C_{\ell_{j}}\,\int\langle\bar{\nabla}\sqrt{\rho_{0}}([\Phi_{J},\Phi_{J^{\perp}}]_{\mathrm{H}}),\,Y_{\ell_{j}}\rangle_{\mathcal{H}_{\mathfrak{m}}}^{2}\,\mathrm{d}\mathfrak{m}^{J^{\perp}}(\Phi_{J^{\perp}})\biggr]$
By combining (5.5) and (5.6) we deduce that
which corresponds to our thesis. ∎
For completeness, below we prove (5.3) and (5.4).
Lemma 5.
$\mathrm{d}\left\|V^{J}_{t}(\overleftarrow{X}^{J}_{t})\right\|_{(\Sigma^{J})^{-1}}^{2}=\biggl(\left\|V^{J}_{t}(\overleftarrow{X}^{J}_{t})\right\|_{(\Sigma^{J})^{-1}}^{2}+\left\|\nabla V^{J}_{t}(\overleftarrow{X}^{J}_{t})\right\|^{2}_{\mathrm{Fr}}\biggr)\mathrm{d}t$
$+2\,V^{J}_{t}(\overleftarrow{X}^{J}_{t})\,\cdot(\Sigma^{J})^{-1}\nabla V^{J}_{t}(\overleftarrow{X}_{t}^{J})\mathrm{d}B_{t}\,$,
Proof.
which implies